# Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v1

@@ II-A RIS-Assisted Wideband THz System Model
where $c$ is the speed of light. Besides, $\mathbf{a}_{\mathrm{T},m}\in\mathbb{C}^{N_{\mathrm{T}}}$ and $\mathbf{a}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}}$ are the normalized array response vectors (ARVs) associated with BS and RIS, respectively, given by
$\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi})=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j2\pi\sin\left(\tilde{\varphi}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
$\mathbf{a}_{\mathrm{R},m}(\tilde{\psi})=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j2\pi\sin\left(\tilde{\psi}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th suncarrier. $d_{\mathrm{BS}}$ and $d^{\mathrm{RIS}}$ denote the antenna spacing of BS and RIS, respectively. Besides, $\mathbf{n}_{t}=\left[0,1,\dots,N_{\mathrm{T}}-1\right]$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]$, respectively. By defining the angle vectors $\tilde{\boldsymbol{\psi}}=\left[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}\right]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be transformed into the virtual angular representation (VAD) form as
$\mathbf{G}[m]=\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\boldsymbol{\varphi}}\right),$ (7)
where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=\left[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})\right]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the AoDs at BS and AoAs at RIS, respectively. $\mathbf{\Upsilon}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-j2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-j2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix. Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and UE-$k$ at the $m$-th SC can be formulated as
@@
where $J_{k}$ represents the number of propagation paths between RIS and UE-$k$, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the AoD at RIS for the $j$-th path, respectively. By defining $\tilde{\boldsymbol{\vartheta}}_{k}=\left[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}\right]$, we transform $\mathbf{h}_{k}[m]$ into the VAD form, given by
@@
where $\boldsymbol{\beta}_{k,m}\!=\!\left[\beta_{k,1}e^{-j2\pi\mu_{k,1}f_{m}},\cdots\!,\beta_{k,J_{k}}e^{-j2\pi\mu_{k,J_{k}}f_{m}}\right]^{\mathsf{T}}$ is the angle-excluded vector for the $k$-th UE. $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the AoDs at RIS. Plugging (7) and (9) into (1), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be written as
$\begin{split}&\mathbf{H}_{k}^{\mathrm{cas}}[m]\\ &=\!\mathrm{Diag}(\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}))\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}})\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\boldsymbol{\beta}_{k,m})\!\bullet\!(\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Upsilon}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}))\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Upsilon}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).\end{split}$ (10)
@@ II-B Beam Split Effect
In RIS-assisted wideband THz systems, beam split effect is essentially caused by the analog beamformer and RIS, which contain the frequency-independent phase shifters [22]. The antenna spacing $d_{\mathrm{BS}}$ and $d_{\mathrm{RIS}}$ are set according to the central frequency $f_{c}$, with half-wavelength spacing, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$ as a typical configuration, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [23]. Substituting the half-wavelength configuration into (6a) and (6b), the ARVs can be rewritten as
$\mathbf{a}_{\mathrm{T},m}(\varphi)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j\pi\eta_{m}\varphi\mathbf{n}_{t}},$ (11a)
$\mathbf{a}_{\mathrm{R},m}(\psi)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j\pi\eta_{m}\psi\mathbf{n}_{r}},$ (11b)
where $\eta_{m}=f_{m}/f_{c}$ represents the relative frequency. Without loss of generality, we define $\varphi=\sin\tilde{\varphi}\in[-1,1]$ and $\psi=\sin\tilde{\psi}\in[-1,1]$ as the physical direction for expression simplicity. In THz band, the physical directions $\varphi$ and $\psi$ split into totally separated spatial directions $\eta_{m}\varphi$ and $\eta_{m}\psi$, which is the so-called beam split effect. Due to the large bandwidth and enormous number of RIS elements, beam split effect occurs at both the BS and RIS, making channel estimation algorithms suffer from severe performance degradation [24].
III Full CSI Estimation by CBS-GAMP
In this section, we propose a novel CBS-GAMP based channel estimation approach where simplified CBS dictionaries are firstly presented to formulate angular domain channel. By employing CBS dictionaries, CBS-GAMP algorithm with EM extensions is implemented alternatively for each angular sub-channel to achieve accurate channel estimation.
@@ III Full CSI Estimation by CBS-GAMP
III-A Angular Domain Sparse Representation
In this subsection, we transformed the THz channel to the angular domain sparse representation by utilizing the overcomplete CBS dictionaries. Specifically, the CBS dictionaries of the BS and RIS directions are formulated as
@@ III-A Angular Domain Sparse Representation
where $Q_{\mathrm{T}}\gg N_{\mathrm{T}}$ and $Q_{\mathrm{R}}\gg N_{\mathrm{R}}$ are the quantization level of the BS and RIS directions, respectively. Besides, $\bar{\vartheta}_{q}=\delta q-\frac{Q+1}{Q}$ with $Q\in\{Q_{\mathrm{T}},Q_{\mathrm{R}}\}$ denotes the spatial direction sample, $\delta=2/Q$ is the interval between two adjacent samples. Substituting the ARMs in (10) by the CBS dictionaries and extending the angle-excluded matrix/vector, the sparse angular domain representation of the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ is given by
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Upsilon}}_{m}\right)\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (13)
where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary, $\bar{\mathbf{\Upsilon}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ none zero entries corresponding to the angle-excluded matrix of BS and the angle-excluded vector of RIS, respectively. According to the operation of transposed Khatri-Rao product and the uniform sampling interval of the CBS dictionary, the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ contains a considerable amount of duplicated columns. For narrowband channel model, the coupled dictionary contains only $Q_{\mathrm{R}}$ distinct columns, which are exactly the first $Q_{\mathrm{R}}$ columns [6]. However, for wideband channel model with beam split effect, due to the frequency-dependent nature of ARVs, this simplification is no longer suitable. Note that sequence $\bar{\vartheta}_{q_{1}}=\delta q_{1}-\frac{Q+1}{Q},\forall q_{1}=1,\dots,Q_{\mathrm{R}}$ corresponds one-to-one with $Q_{\mathrm{R}}$ ARVs in dictionary $\mathbf{C}_{\mathrm{R},m}$, the sequence $-\bar{\vartheta}_{q_{2}},\forall q_{2}=1,\dots,Q_{\mathrm{R}}$ can also establish such correspondence with $\mathbf{C}_{\mathrm{R},m}^{*}$. Therefore, it can be observed that the column vectors of $\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}$ have the form of $\mathbf{a}_{\mathrm{R},m}(\bar{\vartheta}_{q_{1}}-\bar{\vartheta}_{q_{2}}),\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. The $j$-th column of the coupled dictionary $\tilde{{\xi}}_{\mathrm{R},m,j}$ can be expressed as
$\tilde{\xi}_{\mathrm{R},m,j}=\mathbf{a}_{\mathrm{R},m}(\delta\left(q_{1}-q_{2}\right)),$ (14)
where $j=\left(q_{2}-1\right)Q_{\mathrm{R}}+q_{1},\forall q_{1},q_{2}=1,\dots,Q_{\mathrm{R}}$. Based on the fact that $q_{1}-q_{2}\in\left[1-Q_{\mathrm{R}},Q_{\mathrm{R}}-1\right]$, the number of distinct columns of coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{\mathrm{B}}\triangleq 2Q_{\mathrm{R}}-1$. By employing the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, one can easily construct the coupled CBS dictionary as

### v2

@@ II-A Signal and Channel Model
where $f_{c}$ is the central carrier frequency. Besides, $\mathbf{a}_{\mathrm{T},m}\in\mathbb{C}^{N_{\mathrm{T}}}$ and $\mathbf{a}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}}$ are the normalized array response vectors (ARVs) associated with BS and RIS, respectively, given by
$\mathbf{a}_{\mathrm{T},m}(\tilde{u})\triangleq\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j2\pi\sin\left(\tilde{u}\right)\mathbf{n}_{t}d_{\mathrm{BS}}/\lambda_{m}},$ (6a)
$\mathbf{a}_{\mathrm{R},m}(\tilde{v})\triangleq\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j2\pi\sin\left(\tilde{v}\right)\mathbf{n}_{r}d^{\mathrm{RIS}}/\lambda_{m}},$ (6b)
where $\lambda_{m}=c/f_{m}$ is the wavelength of the $m$-th suncarrier. $d_{\mathrm{BS}}$ and $d^{\mathrm{RIS}}$ denote the antenna spacing of BS and RIS, respectively. Besides, $\mathbf{n}_{t}=[0,1,\dots,N_{\mathrm{T}}-1]$ and $\mathbf{n}_{r}=\left[0,1,\dots,N_{\mathrm{R}}-1\right]$, respectively. By defining the angle vectors $\tilde{\boldsymbol{\psi}}=[\tilde{\psi}_{1},\cdots,\tilde{\psi}_{L}]$ and $\tilde{\boldsymbol{\varphi}}=\left[\tilde{\varphi}_{1},\cdots,\tilde{\varphi}_{L}\right]$, $\mathbf{G}[m]$ can be transformed into the virtual angular representation (VAD) form as
$\mathbf{G}[m]=\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}\left(\tilde{\boldsymbol{\varphi}}\right),$ (7)
where $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\psi}_{L})]$ and $\mathbf{A}_{\mathrm{T},m}\left(\tilde{\boldsymbol{\varphi}}\right)=\left[\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{1}),\cdots,\mathbf{a}_{\mathrm{T},m}(\tilde{\varphi}_{L})\right]$ denote the array response matrix (ARM) of the AoDs at BS and AoAs at RIS, respectively. $\mathbf{\Sigma}_{m}=\mathrm{Diag}\left(\alpha_{1}e^{-j2\pi\tau_{1}f_{m}},\cdots,\alpha_{L}e^{-j2\pi\tau_{L}f_{m}}\right)$ is the angle-excluded matrix. Similarly, the channel $\mathbf{h}_{k}[m]\in\mathbb{C}^{1\times N_{\mathrm{R}}}$ between RIS and UE-$k$ at the $m$-th SC can be formulated as
@@
where $J_{k}$ represents the number of propagation paths between RIS and UE-$k$, $\beta_{k,j}$, $\mu_{k,j}$ and $\tilde{\vartheta}_{k,j}$ represent the complex gain, the time delay and the AoD at RIS for the $j$-th path, respectively. By defining $\tilde{\boldsymbol{\vartheta}}_{k}=[\tilde{\vartheta}_{k,1},\cdots,\tilde{\vartheta}_{k,J_{k}}]$, we transform $\mathbf{h}_{k}[m]$ into the VAD form as
@@
where $\boldsymbol{\beta}_{k,m}\!=\!\left[\beta_{k,1}e^{-j2\pi\mu_{k,1}f_{m}},\cdots\!,\beta_{k,J_{k}}e^{-j2\pi\mu_{k,J_{k}}f_{m}}\right]^{\mathsf{T}}$ is the angle-excluded vector for the $k$-th UE. $\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\vartheta}}_{k})=[\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,1}),\cdots,\mathbf{a}_{\mathrm{R},m}(\tilde{\vartheta}_{k,J_{k}})]$ denotes the ARM of the AoDs at RIS. Substituting (7) and (9) into (3), the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ can be rewritten as
$\begin{split}&\mathbf{H}_{k}^{\mathrm{cas}}[m]\\ &=\!\mathrm{Diag}(\boldsymbol{\beta}_{k,m}^{\mathsf{T}}\mathbf{A}_{\mathrm{R},m}^{\mathsf{H}}(\tilde{\boldsymbol{\vartheta}}_{k}))\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}})\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\boldsymbol{\beta}_{k,m})\!\bullet\!(\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}})\mathbf{\Sigma}_{m}\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}))\\ &=(\mathbf{A}_{\mathrm{R},m}^{*}(\tilde{\boldsymbol{\vartheta}}_{k})\!\bullet\!\mathbf{A}_{\mathrm{R},m}(\tilde{\boldsymbol{\psi}}))(\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m})\mathbf{A}_{\mathrm{T},m}^{\mathsf{H}}(\tilde{\boldsymbol{\varphi}}).\end{split}$ (10)
@@ II-B Beam Split Effect
According to (4) and (8), THz channel varies across different SCs due to the beam split effect [^1] The beam split effect can been seen as a severe version of beam squint effect in mmWave band [27].. This phenomenon is caused by the ultra-wide bandwidth and extensive antenna arrays and reflecting elements in RIS-assisted wideband THz systems. Beams generated by the frequency-independent analog beamforming vector and then reflected by the frequency-independent RIS phase shift matrix would split into totally distinct spatial directions. Typically, antenna spacings are set to half the wavelength at the central frequency $f_{c}$, i.e., $d_{\mathrm{BS}}=d_{\mathrm{RIS}}=\lambda_{c}/2$, where $\lambda_{c}=c/f_{c}$ is the wavelength of the central frequency [28]. For simplicity, we define the variables $u\triangleq\sin\tilde{u}\in[-1,1]$ and $v\triangleq\sin\tilde{v}\in[-1,1]$ to represent the physical directions. Substituting the half-wavelength configuration into (6a) and (6b), the ARVs can be rewritten as
$\mathbf{a}_{\mathrm{T},m}(u)=\frac{1}{\sqrt{N_{\mathrm{T}}}}e^{-j\pi\eta_{m}u\mathbf{n}_{t}},$ (11a)
$\mathbf{a}_{\mathrm{R},m}(v)=\frac{1}{\sqrt{N_{\mathrm{R}}}}e^{-j\pi\eta_{m}v\mathbf{n}_{r}},$ (11b)
where $\eta_{m}\triangleq f_{m}/f_{c}$ represents the relative frequency. As shown in (11), the spatial directions can be represented by
$u_{m}=\eta_{m}u,$ (12a)
$v_{m}=\eta_{m}v.$ (12b)
In the lower frequency bands such as mmWave, the subcarrier frequencies are relatively closer, which causes slight beam squint with $\eta_{m}\approx 1$ which means that the ARVs at all SC frequencies are almost identical[28, 20]. However, due to the ultra large bandwidth in THz band, $\eta_{m}$ deviates significantly from one. Therefore, the difference between physical directions and spatial directions becomes significant and cannot be ignored.
In wideband RIS-assisted THz systems, beam split effect occurs at both BS and RIS, which leads to severe performance degradation in channel estimation algorithms [27]. Therefore, it is crucial to develop an effective channel estimation scheme to acquire accurate CSI under beam split effect.
III Channel Estimation
III-A Channel Estimation Protocol
Traditional channel estimation algorithms require significant redesign of the codebook, as conventional dictionaries fail to cover additional angle information introduced by beam split effect [17]. This additional angle information will complicates the design of dictionaries. Furthermore, the beam split effect disperses the channel for each subcarrier into different spatial directions, which requires well-designed channel estimation algorithms to operate on each subcarrier individually, which can significantly increase computational complexity.
To address these challenges, we propose a novel low-complexity channel estimation protocol, which is divided into two main phases, as depicted in Fig. 1. The first phase, which is the most challenging aspect of the protocol, mainly focuses on estimating the full CSI for two selected subcarriers, which includes estimation of cascaded channel, the AoDs at BS, and the coupled angles at RIS. Due to its complexity, we further divide this phase into two steps. In Step 1, we employ sparse recovery algorithms to estimate the cascaded channel for the two subcarriers. In Step 2, we focus on heuristic algorithms for common angles estimation. Given the full CSI estimation from the first phase, we then employ a simple LS algorithm to estimate the channels for the remaining subcarriers.
III-B Cascaded Channel Estimation
In this subsection, we focus on the estimation of cascaded channel for one specific subcarrier. In the following steps, we first present the angular domain sparse representation in RIS-assisted THz channels. Then, we introduce the simplified dictionary, i.e., CBS dictionary. Based on this simplified dictionary, we reformulate the channel estimation problem as a sparse recovery problem. Finally, we provide a detailed analysis of existing sparse recovery algorithms.
III-B1 Angular Domain Sparse Representation
@@
The sparse angular domain representation of the cascaded channel $\mathbf{H}_{k}^{\mathrm{cas}}[m]$ is given by
$\mathbf{H}_{k}^{\mathrm{cas}}[m]=\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\left(\bar{\boldsymbol{\beta}}_{k,m}\otimes\bar{\mathbf{\Sigma}}_{m}\right)\mathbf{C}_{\mathrm{T},m}^{\mathsf{H}},$ (13)
where $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}\triangleq\mathbf{C}_{\mathrm{R},m}^{*}\bullet\mathbf{C}_{\mathrm{R},m}\in\mathbb{C}^{N_{\mathrm{R}}\times Q_{\mathrm{R}}^{2}}$ is the coupled dictionary, $\bar{\mathbf{\Sigma}}_{m}\in\mathbb{C}^{Q_{\mathrm{R}}\times Q_{\mathrm{T}}}$ and $\bar{\boldsymbol{\beta}}_{k,m}\in\mathbb{C}^{Q_{\mathrm{R}}}$ are the sparse formulation with $L$ and $J_{k}$ none zero entries corresponding to the angle-excluded matrix of BS-RIS channel and the angle-excluded vector of RIS-UE channel, respectively. Additionally, the overcomplete dictionaries of the BS and RIS directions can be expressed as
@@
where $Q_{\mathrm{T}}\gg N_{\mathrm{T}}$ and $Q_{\mathrm{R}}\gg N_{\mathrm{R}}$ are the quantization level of the BS and RIS directions, respectively. Moreover, $\bar{\vartheta}_{q}=\delta\cdot q-(Q+1)/{Q}$ with $Q\in\{Q_{\mathrm{T}},Q_{\mathrm{R}}\}$ denotes the spatial direction sample, $\delta=2/Q$ is the interval between two adjacent samples.
III-B2 Simplified Dictionary
Note that the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ in (13) is derived from the transposed Khatri-Rao product, which contains a significant number of redundant columns due to the fixed sampling intervals. For narrowband channel model, this will result in only $Q_{\mathrm{R}}$ unique columns which are exactly the first $Q_{\mathrm{R}}$ columns [9]. However, for wideband THz channel with beam split effect, this simplification is no longer suitable due to the frequency-dependent nature of the ARVs.
Proposition 1.
The number of unique columns of dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$ is $Q_{B}=2Q_{R}-1$, and the coupled dictionary can be simplified by including only the first and the last $Q_{\mathrm{R}}$ columns of the coupled dictionary $\tilde{\mathbf{\Xi}}_{\mathrm{R},m}$, which can be given as